In 5 years from now shyam will be thrice as old as he was 9 years old.Find his present age.
Answer by mananth(16946)

shyam is x years now
after 5 years he will be x+5 years
9 years ago he was x-9 years
x+5=3(x-9)
x+5=3x-27
2x=32
x=16
Shyam is 16 years now